# Measurement Error and Counterfactuals in Quantitative Trade and Spatial Models

### v4

@@ 4 Widely Applicable Default Empirical Bayes Approach
This section proposes a default empirical Bayes (EB) approach that can be applied in many settings. It also discusses the toolkit that accompanies the paper.
@@ 4.1 Default Prior and Measurement Error Model
Often it will be clear what a sensible prior and measurement error model are, for example a Dirichlet prior when observing migration shares. For when this is not the case, in this section I provide a widely applicable default approach for quantifying uncertainty about a counterfactual prediction of interest. This default approach can be applied out-of-the-box to many quantitative trade and spatial models, but can also easily be adapted to other settings. It recommends default choices for the prior distribution and measurement error model, and discusses how to calibrate both based on observed data. [^9] Rather than estimating the parameters of the prior distribution for the true underlying data, which corresponds to an empirical Bayes approach, one could alternatively specify prior distributions for these parameters, which corresponds to a hierarchical Bayes approach.
Concretely, consider the setting where we can write $\gamma=g\left(\left\{F_{ij}\right\},\theta\right)$, for $\left\{F_{ij}\right\}$ a set of non-negative flows between locations. Assume we have access to an estimator $\tilde{\theta}\left(\left\{F_{ij}\right\}\right)$ with estimated sampling variance $\tilde{\Sigma}\left(\left\{F_{ij}\right\}\right)$. This setup is commonplace in quantitative trade and spatial models (Costinot and Rodríguez-Clare, 2014; Redding and Rossi-Hansberg, 2017; Proost and Thisse, 2019).
I assume that both the prior distributions on the true flows and the measurement errors are mixtures of a point mass at zero and a log-normal distribution, a so-called spike-and-slab distribution (Mitchell and Beauchamp, 1988). The point mass at zero is necessary because in both trade and spatial applications bilateral flows of zeros are common, particularly when considering more granular data (Helpman, Melitz, and Rubinstein, 2008; Dingel and Tintelnot, 2020). This prior and measurement error model imply that the posterior distribution of the true flows given the noisy flows will also be a spike-and-slab distribution. This mixture model is fairly flexible and the conjugacy is needed for computational speed. Furthermore, I assume that the prior mean exhibits a gravity relationship, for which there is strong empirical evidence (Head and Mayer, 2014; Allen and Arkolakis, 2018). [^10] One can easily enrich this gravity prior by adding other “distance” variables such as differences in income or productivity, or by adding dummies that indicate similarity such as contiguity or a common language, see for example Silva and Tenreyro (2006). I experimented with this but the results do not change much. This is summarized in the following assumption:
@@
$\begin{cases}\begin{array}[]{c}\mathrm{true\ zeros:}\\ \mathrm{spurious\ zeros:}\\ \mathrm{prior:}\\ \\ \mathrm{likelihood:}\end{array}&\begin{array}[]{c}P_{ij}\sim\mathrm{Bern}\left(p_{ij}\right)\\ B_{ij}\sim\mathrm{Bern}\left(b_{ij}\right)\\ F_{ij}\sim P_{ij}\cdot\delta_{0}+\left(1-P_{ij}\right)\cdot e^{\mathcal{N}\left(\mu_{ij},s_{ij}^{2}\right)}\\ \mu_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}\\ \tilde{F}_{ij}|F_{ij}\sim\delta_{0}\cdot\mathbb{I}\left\{F_{ij}=0\right\}+\left[B_{ij}\cdot\delta_{0}+\left(1-B_{ij}\right)\cdot e^{\mathcal{N}\left(\log F_{ij},\varsigma_{ij}^{2}\right)}\right]\cdot\mathbb{I}\left\{F_{ij}>0\right\},\end{array}\end{cases}$
for $i,j=1,...,n$, where $\mathrm{dist}_{ij}$ denotes the distance between locations $i$ and $j$, $\alpha_{i}^{\mathrm{orig}}$ is an origin fixed effect and $\alpha_{j}^{\mathrm{dest}}$ is a destination fixed effect.
The probability that a bilateral trade flow is truly zero is denoted by $p_{ij}$, and a true zero flow is assumed to always result in an observed zero. The probability of a spurious zero—that is, an observed zero despite a non-zero underlying true flow—is denoted by $b_{ij}$. The prior means and variances are denoted by $\left\{\mu_{ij}\right\}$ and $\left\{s_{ij}^{2}\right\}$, respectively. The flow-specific measurement error variances are denoted by $\left\{\varsigma_{ij}^{2}\right\}$.
Gather the parameters in $\vartheta=\left(\left\{p_{ij}\right\},\left\{b_{ij}\right\},\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},\left\{s_{ij}^{2}\right\},\left\{\varsigma_{ij}^{2}\right\}\right)$. It follows that the posterior distribution for the true flow between location $i$ and $j$, $F_{ij}$, given its noisy version, $\tilde{F}_{ij}$ is given by
@@
for $i,j=1,...,n$, where $Q_{ij}\sim\mathrm{Bern}\left(\frac{p_{ij}}{p_{ij}+b_{ij}\left(1-p_{ij}\right)}\right)$.
Conditional on being able to calibrate the parameters $\vartheta$, one can quantify uncertainty about $\gamma$ by finding the interval $\mathcal{C}^{1}$ as described in Section 3.3. Then, a default procedure for quantifying uncertainty about $\gamma$ is summarized in Algorithm 3.
@@
    Input: noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated prior and likelihood parameters $\tilde{\vartheta}$, structural parameter mapping $\left\{F_{ij}\right\}\mapsto\tilde{\theta},\tilde{\Sigma}$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
@@
    For $b=1,...,B$,
@@
      For $i,j=1,...,n$, draw $F_{ij,b}$ from the posterior distribution $F_{ij}|\tilde{F}_{ij},\tilde{\vartheta}$.
@@
      Sample
      $\theta_{b}\sim\mathcal{N}\left(\tilde{\theta}\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n}\right),\tilde{\Sigma}\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n}\right)\right).$
    3. (c)
      Compute $\gamma_{b}=g\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n},\theta_{b}\right)$.
  3. 3.
@@
with the probability density function of a standard normal distribution. To further check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity model and, following Allen and Arkolakis (2018), plot the log flows against the log distance for positive flows, after partitioning out the origin and destination fixed effects. In Appendices G and H I perform both these checks for my applications.
@@
Specifying a prior and measurement error model is difficult and one might be worried about misspecification. For the normal-normal model, we can use prior density-ratio classes to find worst-case bounds on posterior quantiles over a neighborhood that contains distributions that are not too far away from the assumed normal distribution for the prior and measurement error model. It turns out that incorporating uncertainty around the prior and measurement error model amounts to reporting slightly wider quantiles. The details can be found in Appendix D.
Remark 3.
In this default approach, one might worry about attenuation bias when plugging in the shrunk data $\left\{F_{ij,b}\right\}_{i,j=1}^{n}$ into the estimator $\tilde{\theta}$. Such bias would indeed arise if the data were shrunk toward zero or another constant (Chen, Gu, and Kwon, 2025). However, the default approach instead shrinks toward an economically motivated gravity prior, whose fitted values are expected to serve as a reasonable proxy for the true flows. A simulation exercise illustrating this point, based on the running example, is provided in Appendix E.
4.2 Empirical Bayes Step: Calibrating $\vartheta$
The parameters in $\vartheta$ need to be calibrated. In consider two cases.
@@ 4.2.1 Baseline Case with Domain Knowledge
In the baseline case I restrict the measurement error variance and prior variance to be constant across flows so that $\varsigma_{ij}^{2}=\varsigma^{2}$ and $s_{ij}^{2}=s^{2}$ for all $i,j=1,...,n$. Furthermore, I require knowledge of the common measurement error variance $\varsigma^{2}$ and of the Bernoulli parameters $\left\{p_{ij}\right\}$ and $\left\{b_{ij}\right\}$. [^11] In the absence of a prior on the measurement error variance, one could adopt a sensitivity analysis approach by varying the variance to determine the minimum level of measurement error that would overturn the counterfactual conclusion. It then remains to estimate $\left(\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},s^{2}\right)$. Towards this, we can combine the equations in Assumption 2 to find
@@
Obtaining estimators for these prior means and variances is what Walters (2024) calls the deconvolution step.
@@ 4.2.2 Mirror Trade Data
When the non-negative bilateral flows correspond to trade flows between countries, I use the mirror trade dataset from Linsi, Burgoon, and Mügge (2023) to calibrate $\vartheta$. This dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow. The details for the calibration can be found in Appendix F. I first calibrate the probabilities of true zeros $\left\{p_{ij}\right\}$ and the probabilities of spurious zeros $\left\{b_{ij}\right\}$ by noting that for each bilateral trade flow we can use the time variation to identify the probabilities of observing a certain number of zeros. I then leverage the model structure to calibrate the measurement error variances $\left\{\varsigma_{ij}^{2}\right\}$. Lastly, I calibrate the prior parameters, which are period-specific in this case, using a similar approach as for the baseline case with domain knowledge.
@@ 4.3 Toolkit
Accompanying the paper, I provide an easy-to-use toolkit that consists of three programs. [^12] The toolkit is written in MATLAB and can be found on my website, https://sandersbas.github.io/. A version in R is available upon request. The first program implements the high-level approach in Algorithm 1. It takes $\left(B,\tilde{D},\pi^{ME},\tilde{\theta},\tilde{\Sigma},g\right)$ as inputs and outputs posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$. The second program implements the default approach in Algorithm 3. It takes as inputs $\left(B,\left\{\tilde{F}_{ij}\right\},\tilde{\vartheta},\tilde{\theta},\tilde{\Sigma},g\right)$ and again outputs posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$. The third program, which can serve as an input to the second, uses the mirror trade dataset of Linsi, Burgoon, and Mügge (2023) and allows the researcher to choose countries and years for which they want to estimate the parameters of the prior and measurement error model. This is summarized in Algorithm 4.
@@
    • Input: number of draws $B$, data $\tilde{D}$, functions $\tilde{D}\mapsto D_{b}$, $D\mapsto\left(\tilde{\theta},\tilde{\Sigma}\right)$, $\left(D,\theta\right)\mapsto\gamma$.
@@
    Program 2: Default Approach
    • Input: number of draws $B$, noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated parameters $\tilde{\vartheta}$, functions, $\left\{F_{ij}\right\}\mapsto\left(\tilde{\theta},\tilde{\Sigma}\right)$, ($\left(\left\{F_{ij}\right\},\theta\right)\mapsto\gamma$.
@@
    • Output: noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated parameters $\tilde{\vartheta}$, adjusted R-squared of the gravity model for the last year in $\mathcal{T}$, plot of log flows against log distance for positive flows, after partitioning out the origin and destination fixed effects as per Remark 1.

### v5

@@ 4 Widely Applicable Default Approach
This section proposes a default empirical Bayes approach that can be applied in many settings, as it is both economically reasonable for many quantitative trade and spatial models and computationally convenient. It also discusses the toolkit that accompanies the paper.
@@ 4.1 Default Prior and Measurement Error Model
In many applications, the support of the true data and the sampling process naturally suggest both a measurement error model and a prior. For example, when observing migration shares constructed from count data, a multinomial measurement error model arises naturally from sampling variation. Since the true shares are non-negative and must sum to one, a Dirichlet prior provides a coherent and tractable specification.
For settings where such natural choices are not available, this section proposes a widely applicable default approach for quantifying uncertainty about the counterfactual prediction of interest. This default approach can be applied out-of-the-box to many quantitative trade and spatial models, but can also be easily adapted to other settings. It recommends default choices for the prior distribution and measurement error model, and discusses how to calibrate both based on observed data.
Concretely, consider the setting where we can write $\gamma=g\left(\left\{F_{ij}\right\},\theta\right)$, for $\left\{F_{ij}\right\}$ a set of non-negative flows between locations. This setup is commonplace in quantitative trade and spatial models (costinot2014trade; redding2017quantitative; proost2019can). I assume that both the prior distributions on the true flows and the measurement errors are mixtures of a point mass at zero and a log-normal distribution, a so-called spike-and-slab distribution (mitchell1988bayesian). The point mass at zero is necessary because in both trade and spatial applications bilateral flows of zeros are common, particularly when considering more granular data (helpman2008estimating; DingelTintelnot:2025). This prior and measurement error model imply that the posterior distribution of the true flows given the noisy flows will also be a spike-and-slab distribution. This mixture model is fairly flexible, and conjugacy ensures computational tractability. Furthermore, I assume that the prior mean exhibits a gravity relationship, for which there is strong empirical evidence (head2014gravity; allen201813). [^5] One can easily enrich this gravity prior by adding other “distance” variables such as differences in income or productivity, or by adding dummies that indicate similarity such as contiguity or a common language, see for example silva2006log. I experimented with this but the results do not change much. This is summarized in the following assumption:
@@
$\begin{cases}\begin{array}[]{c}\mathrm{true\ zeros:}\\ \mathrm{spurious\ zeros:}\\ \mathrm{prior:}\\ \\ \mathrm{likelihood:}\\ \\ \end{array}&\begin{array}[]{c}P_{ij}\sim\mathrm{Bern}\left(p_{ij}\right)\\ B_{ij}\sim\mathrm{Bern}\left(b_{ij}\right)\\ F_{ij}\sim P_{ij}\cdot\delta_{0}+\left(1-P_{ij}\right)\cdot e^{\mathcal{N}\left(\mu_{ij},s_{ij}^{2}\right)}\\ \mu_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}\\ \tilde{F}_{ij}|F_{ij}\sim\delta_{0}\cdot\mathbb{I}\left\{F_{ij}=0\right\}\\ +\left[B_{ij}\cdot\delta_{0}+\left(1-B_{ij}\right)\cdot e^{\mathcal{N}\left(\log F_{ij},\varsigma_{ij}^{2}\right)}\right]\cdot\mathbb{I}\left\{F_{ij}>0\right\}\end{array},\end{cases}$
for i,j=1,,ni,j=1,...,n, where $\delta_{0}$ denotes the Dirac mass at zero, $\mathrm{dist}_{ij}$ denotes the distance between locations $i$ and $j$, $\alpha_{i}^{\mathrm{orig}}$ is an origin fixed effect and $\alpha_{j}^{\mathrm{dest}}$ is a destination fixed effect.
The probability that a bilateral trade flow is truly zero is denoted by $p_{ij}$, and a true zero flow is assumed to always result in an observed zero. [^6] Assumption 2 implies that both true and spurious zeros occur randomly. Alternatively, one could think about incorporating endogenous zeros using selection mechanisms such as in helpman2008estimating. The probability of a spurious zero—that is, an observed zero despite a non-zero underlying true flow—is denoted by $b_{ij}$. The prior means and variances are denoted by $\left\{\mu_{ij}\right\}$ and $\left\{s_{ij}^{2}\right\}$, respectively. The flow-specific measurement error variances are denoted by $\left\{\varsigma_{ij}^{2}\right\}$.
Gather the hyperparameters in $\vartheta=\left(\left\{p_{ij}\right\},\left\{b_{ij}\right\},\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},\left\{s_{ij}^{2}\right\},\left\{\varsigma_{ij}^{2}\right\}\right)$. It follows that the posterior distribution for the true flow between location $i$ and $j$, $F_{ij}$, given its noisy version, $\tilde{F}_{ij}$ is given by
@@
for i,j=1,,ni,j=1,...,n, where $Q_{ij}\sim\mathrm{Bern}\left(\frac{p_{ij}}{p_{ij}+b_{ij}\left(1-p_{ij}\right)}\right)$. [^7] Note that $F_{ij}$ are drawn independently across $ij$-pairs. As a result, certain adding-up constraints may no longer hold. In principle, this issue could be addressed by drawing from a constrained joint distribution.
Conditional on being able to calibrate the parameters $\vartheta$—the empirical Bayes estimation step—one can quantify uncertainty about $\gamma$ by finding the interval as described in Algorithm 1. Then, a default procedure for quantifying uncertainty about $\gamma$ is summarized in Algorithm 2.
@@
    Input: noisy flows $\left\{\tilde{F}_{ij}\right\}$, structural parameter $\theta$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
@@
    Empirical Bayes estimation step: calibrate $\vartheta$ as outlined in Section 4.2 and denote the estimator by $\tilde{\vartheta}$.
  3. 3.
    For b=1,,Bb=1,...,B,
@@
[…]
@@
      Compute $\gamma_{b}=g\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n},\theta\right)$.
  4. 4.
@@
with the probability density function of a standard normal distribution. To further check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity model and, following allen201813, plot the log flows against the log distance for positive flows, after partitioning out the origin and destination fixed effects. In Appendices D and E I perform both these checks for my applications.
@@
One might be worried about misspecification of the prior and measurement error model. For the normal-normal model, we can use prior density-ratio classes to find worst-case bounds on posterior quantiles over a neighborhood that contains distributions that are not too far away from the assumed normal distribution for the prior and measurement error model. It turns out that incorporating uncertainty around the prior and measurement error model amounts to reporting slightly wider quantiles. The details can be found in Appendix F.
4.2 Empirical Bayes Estimation Step: Calibrating $\vartheta$
The hyperparameters in $\vartheta$ need to be calibrated. I consider two cases.
@@ 4.2.1 Baseline Case with Domain Knowledge
In the baseline case I restrict the measurement error variances and prior variances to be constant across flows so that $\varsigma_{ij}^{2}=\varsigma^{2}$ and $s_{ij}^{2}=s^{2}$ for all i,j=1,,ni,j=1,...,n. Furthermore, I require knowledge of the common measurement error variance $\varsigma^{2}$ and of the Bernoulli parameters $\left\{p_{ij}\right\}$ and $\left\{b_{ij}\right\}$. [^8] In the absence of a prior on the measurement error variance, one could adopt a sensitivity analysis approach by plotting intervals while varying the variance. Alternatively, one could determine the minimum level of measurement error that would overturn the counterfactual conclusion. It then remains to estimate $\left(\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},s^{2}\right)$. Towards this, we can combine the equations in Assumption 2 to find
@@
Obtaining estimators for these prior means and variances is what walters2024empirical calls the deconvolution step.
@@ 4.2.2 Mirror Trade Data
When the non-negative bilateral flows correspond to trade flows between countries, I use the mirror trade dataset from linsi2023problem to calibrate $\vartheta$. This dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. linsi2023problem shows that there are so-called mirror discrepancies in bilateral trade flows between almost all countries. This means that, for instance, while the value that Germany reports it imported from France and the value that France reports it exported to Germany should be the same, in practice they are often different. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow. The key identifying assumptions are that the flow-specific probabilities of true zeros, the flow-specific probabilities of spurious zeros, and the flow-specific measurement error variances are constant over time.
The details for the calibration can be found in Appendix G. I first calibrate the probabilities of true zeros $\left\{p_{ij}\right\}$ and the probabilities of spurious zeros $\left\{b_{ij}\right\}$ by noting that for each bilateral trade flow we can use the time variation to identify the probabilities of observing a certain number of zeros. I then leverage the model structure to calibrate the measurement error variances $\left\{\varsigma_{ij}^{2}\right\}$. Lastly, I calibrate the prior parameters, using a similar approach as for the baseline case with domain knowledge. To leverage country information and the fact that importers and exporters can differ in their reliability, I shrink the measurement error and prior variances using country-origin and country-destination fixed effects.
@@ 4.3 Toolkit
Accompanying the paper, I provide an easy-to-use toolkit that consists of three programs. [^9] The toolkit is written in MATLAB and can be found on my website, https://sandersbas.github.io/. A version in R is available upon request. The first program implements the high-level approach in Algorithm 1. It takes as inputs $\left(B,\theta,\tilde{D},\pi^{\mathrm{post}}\left(D|\tilde{D},\tilde{\vartheta}\right),g\right)$ and outputs posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$. The second program implements the default approach in Algorithm 2. It takes as inputs $\left(B,\theta,\left\{\tilde{F}_{ij}\right\},\tilde{\vartheta},g\right)$ and again outputs posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$. The third program, which can serve as an input to the second, uses the mirror trade dataset of linsi2023problem and allows the researcher to choose countries and years for which they want to estimate the hyperparameters of the prior and measurement error model. This is summarized in Algorithm 3.
@@
    • Input: number of draws $B$, structural parameter $\theta$, data $\tilde{D}$, functions $\tilde{D}\mapsto D_{b}$, $\left(D,\theta\right)\mapsto\gamma$.
@@
    Program 2: Default approach.
    • Input: number of draws $B$, structural parameter $\theta$, noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated hyperparameters $\tilde{\vartheta}$, function $\left(\left\{F_{ij}\right\},\theta\right)\mapsto\gamma$.
@@
    • Output (can serve as input to Program 2): noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated hyperparameters $\tilde{\vartheta}$, adjusted R-squared of the gravity model for the last year in $\mathcal{T}$, plot of log flows against log distance for positive flows, after partitioning out the origin and destination fixed effects as per Remark 1.
5 Prototypical Example: Armington Model
This section illustrates the proposed procedure using the Armington model (armington1969theory), a canonical workhorse model in international trade, as outlined, for example, in costinot2014trade.
5.1 Model and Counterfactual Question of Interest
Countries are indexed by i,j=1,,ni,j=1,...,n, and with CES preferences and perfect competition, it follows that the relevant gravity equations and budget constraints are:
$E_{i}$ $=\left(1+\kappa_{i}\right)Y_{i},\qquad i=1,...,n.$ (11)
Here, $F_{ij}$ denotes the trade flow from country $i$ to $j$, and $Y_{i}=\sum_{\ell=1}^{n}F_{i\ell}$, $E_{i}=\sum_{k=1}^{n}F_{ki}$ and $\kappa_{i}=\left(E_{i}-Y_{i}\right)/Y_{i}$ denote country $i$’s total income, total expenditure and the ratio of the trade deficit to income, respectively. Furthermore, $\tau_{ij}$ denotes the iceberg trade cost between country $i$ and $j$, which means that in order to sell one unit of a good in country $j$, country $i$ must ship $\tau_{ij}\geq 1$ units, with $\tau_{ii}=1$. Lastly, $\varepsilon>0$ is the trade elasticity and $\left\{\chi_{ij}\right\}$ are idiosyncratic preferences.
Now, say we are interested in the counterfactual where we change the trade costs $\left\{\tau_{ij}\right\}$ proportionally by $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, holding the trade elasticity $\varepsilon$, the idiosyncratic preferences $\left\{\chi_{ij}\right\}$ and the trade imbalance variables $\left\{\kappa_{i}\right\}$ constant. In Appendix H.1 I show that we can then solve for the corresponding proportional changes in income, $\left\{Y_{i}^{\mathrm{cf,prop}}\right\}$, using
where $\lambda_{ij}=F_{ij}/E_{j}$ denotes the expenditure share that country $j$ spends on goods from country $i$. By Walras’ Law, the proportional changes in income are only pinned down up to a multiplicative constant. Subsequently, following costinot2014trade, we can exactly solve for proportional changes in expenditure shares and welfare (real consumption) levels:
$\lambda_{ij}^{\mathrm{cf,prop}}$ $=\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}{\sum_{k}\lambda_{kj}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}},\qquad i,j=1,...,n$
$D$ $=\left\{F_{ij}\right\}$
$\theta$ $=\varepsilon.$
The specific counterfactual question I consider is a 10% increase in all bilateral trade costs between 76 countries, so that $\tau_{ij}^{\mathrm{cf,prop}}=1+0.1\cdot\mathbb{I}\left\{i\neq j\right\}$ for i,j=1,,ni,j=1,...,n. I focus on the proportional percentage changes in welfare in the Central African Republic, the Netherlands, Sweden and the United States. It follows that, fixing $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, we have
$\gamma_{q}=100\cdot\left(W_{q}^{\mathrm{cf,prop}}-1\right)\equiv g_{q}\left(\left\{F_{ij}\right\},\varepsilon\right),$ (12)
for each $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$.
5.2 Measurement Error Model and Prior
For the Armington model, I will consider measurement error in trade flows $\left\{F_{ij}\right\}$. Hence, instead of the true trade flows we observe noisy trade flows $\left\{\tilde{F}_{ij}\right\}$, which in turn lead to noisy counterfactual predictions $\tilde{\gamma}_{q}$ for $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$. Our goal is to quantify the uncertainty that arises from measurement error, and report an accompanying point estimator that aims to answer the question what a Bayesian believes the counterfactual predictions would have been in the absence of measurement error.
[…]
5.3 Results
We can see the impact of measurement error in Table 1 and Figure 1. In Table 1 I compare the standard point estimates based on noisy flows and the posterior median estimates, and report the intervals obtained using Algorithm 2. In Figure 1 I plot the standard point estimates and the smoothed estimated posterior distributions.
We observe that for the Central African Republic and the Netherlands there is a considerable difference between the point estimate and the median of the posterior distribution, causing the point estimate to lie outside the credible set. For Sweden and the United States there is less of a discrepancy. These plots illustrate that the proposed approach automatically incorporates bias induced by measurement error, and that this bias can be both negative and positive. In Appendix H.2, I show the results for all 76 countries in the sample.
[…]
$\gamma_{\mathrm{CAF}}$ -1.09 -0.26 [-0.42, -0.14]
$\gamma_{\mathrm{NLD}}$ -5.15 -6.57 [-7.10, -6.04]
$\gamma_{\mathrm{SWE}}$ -3.25 -3.51 [-3.79, -3.25]
$\gamma_{\mathrm{USA}}$ -1.07 -1.01 [-1.27, -0.56]
[…]